Strings join two object, whose masses are 5 kg and 10 kg. The 5kg object is pulled upward with a force of 120N. Find the tension on the string between the objects.
1
Expert's answer
2020-12-15T11:32:58-0500

If the 120-N force is applied to the string of 5-kg mass, and the 10-kg mass is attached to the 5-kg one with a string, the tensions will be


"T=120+gM_{10}=\\=120+9.8\cdot10=218\text{ N}."